# Improving Retrieval-Augmented Generation through Multi-Agent Reinforcement Learning

### v1

@@ Abstract.
Retrieval-augmented generation (RAG) is extensively utilized to incorporate external, current knowledge into large language models, thereby minimizing hallucinations. A standard RAG pipeline may comprise several components, such as query rewriting, document retrieval, document filtering, and answer generation. However, these components are typically optimized separately through supervised fine-tuning, which can lead to misalignments between the objectives of individual modules and the overarching aim of generating accurate answers in question-answering (QA) tasks. Although recent efforts have explored reinforcement learning (RL) to optimize specific RAG components, these approaches often focus on overly simplistic pipelines with only two components or do not adequately address the complex interdependencies and collaborative interactions among the modules. To overcome these challenges, we propose treating the RAG pipeline as a multi-agent cooperative task, with each component regarded as an RL agent. Specifically, we present MMOA-RAG, a Multi-Module joint Optimization Algorithm for RAG, which employs multi-agent reinforcement learning to harmonize all agents’ goals towards a unified reward, such as the F1 score of the final answer. Experiments conducted on various QA datasets demonstrate that MMOA-RAG improves the overall pipeline performance and outperforms existing baselines. Furthermore, comprehensive ablation studies validate the contributions of individual components and the adaptability of MMOA-RAG across different RAG components and datasets [^1] The code of MMOA-RAG is on https://github.com/chenyiqun/MMOA-RAG..
Retrieval-Augmented Generation; Multi-Agent Cooperation; Multi-Agent Reinforcement Learning; Multi-Module Joint Learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language generationccs: Information systems Question answering
1. Introduction
@@
Large Language Models (LLMs) have been widely applied to tasks such as question answering (Asai et al., 2023; Khattab et al., 2022), information retrieval (Sun et al., 2023; Chen et al., 2024), various forms of reasoning (Huang and Chang, 2022; Hao et al., 2023), and evaluation (Gong and Mao, 2023; Fu et al., 2023). However, since LLMs cannot promptly update their internal knowledge after pre-training, they remain prone to generating outdated or fabricated responses (Zhao et al., 2023). To address these challenges, Retrieval-Augmented Generation (RAG) enhances the generative capabilities of LLMs by retrieving relevant information from external knowledge sources. Recent RAG systems are often built as complex pipelines comprising multiple interconnected modules  (Gao et al., 2023), including query rewriting (Ma et al., 2023; Jiang et al., 2024), first-stage retrieval (Lewis et al., 2020; Shao et al., 2023), re-ranking (Salemi and Zamani, 2024a, b), document pre-processing (Ke et al., 2024; Li et al., 2024b), and answer generation (Shao et al., 2023; Su et al., 2024).
@@
The complexity of RAG systems makes their optimization particularly challenging. Standard supervised fine-tuning (SFT) optimizes each module independently using human-annotated data. However, this often results in misalignment between the objectives of individual components and the system’s overarching goal of generating high-quality outputs. For example, retrieval modules are frequently trained on human-labeled relevance data to optimize metrics such as nDCG(Järvelin and Kekäläinen, 2002), but this process fails to address the disconnect between document relevance and response quality—documents with high relevance scores do not always contribute to generating accurate answers  (Cuconasu et al., 2024).
@@
To address this issue, existing work on end-to-end optimization for RAG, such as  (Lee et al., 2019; Guu et al., 2020; Lewis et al., 2020; Salemi and Zamani, 2024b) aims to propagate rewards from the final output to intermediate modules using techniques like attention distributions (Izacard and Grave, 2020), generation probability (Lewis et al., 2020; Zamani and Bendersky, 2024), and expectation-maximization (EM) iterations (Singh et al., 2021; Salemi and Zamani, 2024a). However, earlier approaches primarily focus on simplified pipelines with only two components-a retriever and a generator-and fail to provide a generalizable framework for jointly optimizing complex systems with multiple components and richer interdependencies. More recent methods attempt to eliminate the need for module-specific rewards by leveraging algorithms like Direct Preference Optimization (DPO) (Rafailov et al., 2024) and Proximal Policy Optimization (PPO)(Schulman et al., 2017). Nonetheless, these methods still concentrate on optimizing individual RAG modules in isolation, without adequately modeling the collaborative dynamics between interacting components (Li et al., 2024b; Ma et al., 2023; Ke et al., 2024). Effectively capturing interdependencies among multiple modules and jointly optimizing complex RAG architectures remains an open research challenge.
Figure 1. The overall framework of MMOA-RAG.
@@
In this paper, we propose a novel approach called the Multi-Module joint Optimization Algorithm (MMOA-RAG) to enable joint optimization across multiple modules in a RAG system. Our framework treats each intermediate component in the RAG pipeline as an agent and formulates the optimization process as a multi-agent collaborative reinforcement learning (RL) problem, where the agents (i.e., modules) work together to maximize a shared reward for the final outcome.
@@
While MMOA-RAG is flexible with the choices reward function and pipeline design, in this study, we define the final reward as the correctness of the generated response, measured by the F1 score against the ground-truth answer, and apply MMOA-RAG to a RAG pipeline comprising four modules: a query rewriter, a fixed document retriever, a document selector, and an answer generator. To achieve this, we employ the Multi-Agent PPO (Yu et al., 2022) algorithm, enabling collaborative optimization within a fully cooperative setting. In this setting, the optimization objectives of all modules are aligned toward the ultimate goal of generating high-quality answers.
Compared to previous approaches for RAG, MMOA-RAG allows for the end-to-end optimization of complex RAG systems, ensuring that each module’s objectives align with the overarching goal of producing accurate responses. Additionally, unlike recent methods based on DPO (Zhu et al., 2024a, b) or PPO (Ma et al., 2023; Ke et al., 2024), our approach is more flexible for various pipeline designs and excels at fostering collaboration among multiple modules.
@@
To demonstrate the effectiveness of the MMOA-RAG modeling and optimization approach, we conducted experiments on three publicly available QA datasets, HotpotQA (Yang et al., 2018), 2WikiMultihopQA (Ho et al., 2020) and AmbigQA (Min et al., 2020), based on Llama-3-8B-Instruct (Dubey et al., 2024). The experimental results indicate that MMOA-RAG achieves better performance than a series of existing optimization methods for RAG. Additionally, we performed extensive ablation studies to investigate the effectiveness and advantages of jointly optimizing mulitple modules in the RAG system and the generalizability of MMOA-RAG across different RAG pipelines.
@@
  • We employ a multi-agent cooperative reinforcement learning algorithm to jointly optimize a sophisticated RAG system with three modules: a query writer, a document selector, and an answer generation.
@@
  • We conduct extensive experiments to verify and demonstrate the effectiveness of the proposed framework.
2. Related Works
2.1. End-to-end optimization in OpenQA
ORQA (Lee et al., 2019) is an open-domain QA system that learns end-to-end evidence retrieval and answer generation using only question-answer pairs, enabled by pretraining with an Inverse Cloze Task. REALM (Guu et al., 2020) is an end-to-end optimizing framework that enhances language model pre-training with a retrieval-augmented approach. Lewis et al. (2020) introduces Retrieval-Augmented Generation as RAG, a model that combines pre-trained language models with non-parametric memory for improved performance on knowledge-intensive NLP tasks. Izacard and Grave (2020) propose a knowledge distillation method to train retriever models using synthetic labels derived from reader model attention scores. Stochastic RAG (Zamani and Bendersky, 2024) introduces a novel end-to-end optimization framework for RAG through expected utility maximization.
2.2. RAG without parameters update
These methods typically involve designing a novel RAG mechanism to enhance the performance of LLMs on question answering tasks. For example, DSP (Khattab et al., 2022) leverages sophisticated interactions between retrieval and language models to address knowledge-intensive NLP tasks. FLARE (Jiang et al., 2023), an active retrieval augmented generation method, enhances text generation by dynamically retrieving relevant information throughout the process, showing superior performance across various long-form knowledge-intensive tasks. ITER-RETGEN (Shao et al., 2023) is an iterative retrieval-generation synergy method that enhances retrieval-augmented large language models by synergistically combining retrieval and generation in an iterative manner. Search-in-the-Chain (Xu et al., 2024a) is a framework that interactively enhances Large Language Models with search capabilities to improve performance on complex, knowledge-intensive tasks. SELF-RAG (Asai et al., 2023) enhances language model quality and factuality through self-reflective retrieval and generation. DRAGIN (Su et al., 2024) is a dynamic RAG framework that addresses the real-time information needs of LLMs during text generation, enhancing performance on knowledge-intensive tasks. GenGround (Shi et al., 2024) synergizes large language model knowledge with external documents to enhance multi-hop question answering through an iterative process of generating answers and grounding them in evidence. Astute RAG (Wang et al., 2024) is a approach that enhances the robustness of Retrieval-Augmented Generation for Large Language Models by adaptively integrating internal and external knowledge while resolving knowledge conflicts.
2.3. RAG with parameters update
2.3.1. Optimizing RAG with SFT
INFO-RAG (Xu et al., 2024b), an unsupervised training method, enhances the capacity of large language models to integrate and refine information from retrieved texts. LongRAG (Zhao et al., 2024b) introduces a dual-perspective retrieval-augmented generation system to enhance understanding of complex long-context knowledge for improved performance in long-context question answering tasks. In INSTRUCTRAG (Wei et al., [n. d.]), generation accuracy and trustworthiness are enhanced by explicitly denoising retrieved information through self-synthesized rationales, outperforming standard RAG approaches without additional supervision.
2.3.2. Optimizing RAG with RL
Some existing works use PPO (Schulman et al., 2017) algorithm to fine-tune LLMs. In Rewrite-Retrieve-Read framework (Ma et al., 2023), a small language model is trained with reinforcement learning to rewrite queries for RAG. BGM (Ke et al., 2024) proposes a novel bridge mechanism between retrieval model and LLMs and uses PPO to optimize the parameters of the bridge to filter for more helpful documents. SMARTRAG (Gao et al., 2024) optimizes an iterative RAG framework with reward, which includes a decision maker and a policy network. RAG-Star (Jiang et al., 2024) is a reasoning approach that combines Monte Carlo Tree Search (MCTS) to improve the complex reasoning abilities of LLMs.
Some other works use DPO (Rafailov et al., 2024) or similar alignment algorithms to optimize LLMs. A noise-filtering method (Zhu et al., 2024a) is proposed by optimizing mutual information between compressed data and output while minimizing it with the retrieved passage. ATM (Zhu et al., 2024b), an Adversarial Tuning Multi-agent system, enhances the robustness and performance of retrieval-augmented generators in question answering by iteratively tuning against an adversarial attacker agent to better discriminate useful documents and resist fabricated content. SEER (Zhao et al., 2024a) proposes a novel self-aligned evidence extraction learning framework aimed at enhancing RAG performance by optimizing the extraction of high-quality, concise, and relevant evidence. RAG-DDR (Li et al., 2024b) optimizes RAG systems by aligning data preferences between modules through DDR, resulting in enhanced performance on knowledge-intensive tasks.
3. Preliminary
3.1. Modeling RAG as Co-MARL
@@
In this context, we define the tuple $\left\langle\mathcal{G},\mathcal{O},\mathcal{A},\mathcal{R}\right\rangle$, where $\mathcal{G}$ denotes the set of agents in the Co-MARL system, $\mathcal{O}$ represents the observation information available to each agent, $\mathcal{A}$ constitutes the action space accessible to each agent, and $\mathcal{R}$ is the reward shared among all agents. The ultimate aim is to maximize this shared reward, thereby achieving higher evaluation metrics and enhancing the overall performance of the RAG system.
3.2. MAPPO Algorithm
In this paper we utilizes Multi-Agent PPO (MAPPO) (Yu et al., 2022), which is an extension of the PPO algorithm (Schulman et al., 2017) for multi-agent environments, to optimize the policy for each agent in the Co-MARL framework. In fully cooperative settings, unlike PPO, which focuses on single-agent scenarios with individual reward, MAPPO employs a shared global reward to promote cooperation among all agents. Additionally, a further distinction between MAPPO and PPO lies in the input received by their critic models. In PPO, the critic model is limited to the agent’s observation information, whereas MAPPO’s global critic model can access comprehensive global information, which enables the global critic model to more accurately estimate the state-value function.
4. Method
4.1. Overall of MMOA-RAG
@@
  • Query Rewriter reformulates the initial query $q$, which may be too complex or ambiguous to resolve with a single retrieval, into a set of sub-questions denoted as $subq$.
@@
Since the Query Rewriter, Selector, and Generator modules can all be implemented using LLMs, they can be treated as RL agents (Ouyang et al., 2022), enabling parameter updates through reward signals. To optimize computational efficiency, these three modules can share the same LLM. Additionally, given the difficulty of modeling the Retriever module as an RL agent, we consider the Retriever to be part of the environment  [^2] Recent studies in generative IR (see  Li et al. (2024a) for a survey) have explored using generative models for retrieval. But we choose a more traditional dense retrieval model (Izacard et al., 2021) as the first-stage retriever and leave the optimization of the first-stage retriever for future work..
@@
The focus of the MMOA-RAG framework is on the collaborative optimization of multiple modules to align their individual optimization objectives with the ultimate goal of generating high-quality answers. We can use metrics derived from the Generator’s predicted answer $Ans_{\text{predict}}$, such as the F1 score, as a shared reward $R_{\text{shared}}$. Given the fully cooperative nature of the modules in the RAG system, the shared reward $R_{\text{shared}}$ can be used to train all agents, a common approach in existing MARL literature (Rashid et al., 2018; Yu et al., 2022; Chen et al., 2022). Additionally, to ensure training stability and accelerate convergence in the multi-agent system, we design penalty terms $P_{QR}$, $P_{S}$, and $P_{G}$ for each agent. A more detailed explanation will be provided in Section 4.2.
4.2. Detailed Configuration for Each Agent

### v2

@@ Abstract
Retrieval-augmented generation (RAG) is widely utilized to incorporate external knowledge into large language models, thereby enhancing factuality and reducing hallucinations in question-answering (QA) tasks. A standard RAG pipeline consists of several components, such as query rewriting, document retrieval, document filtering, and answer generation. However, these components are typically optimized separately through supervised fine-tuning, which can lead to misalignments between the objectives of individual components and the overarching aim of generating accurate answers. Although recent efforts have explored using reinforcement learning (RL) to optimize specific RAG components, these approaches often focus on simple pipelines with only two components or do not adequately address the complex interdependencies and collaborative interactions among the modules. To overcome these limitations, we propose treating the complex RAG pipeline with multiple components as a multi-agent cooperative task, in which each component can be regarded as an RL agent. Specifically, we present MMOA-RAG [^1] The code of MMOA-RAG is on https://github.com/chenyiqun/MMOA-RAG., Multi-Module joint Optimization Algorithm for RAG, which employs multi-agent reinforcement learning to harmonize all agents’ goals toward a unified reward, such as the F1 score of the final answer. Experiments conducted on various QA benchmarks demonstrate that MMOA-RAG effectively boost the overall performance of the pipeline and outperforms existing baselines. Furthermore, comprehensive ablation studies validate the contributions of individual components and demonstrate MMOA-RAG can be adapted to different RAG pipelines and benchmarks.
1 Introduction
@@
Large Language Models (LLMs) have been widely applied to tasks such as question answering asai2023self ; khattab2022demonstrate , information retrieval sun2023chatgpt ; chen2024tourrank , various forms of reasoning huang2022towards ; hao2023reasoning , and evaluation gong2023coascore ; fu2023gptscore . However, since LLMs cannot promptly update their internal knowledge after pre-training, they are still prone to generating outdated or fabricated responses zhao2023survey . To address these challenges, Retrieval-Augmented Generation (RAG) enhances the generative capabilities of LLMs by retrieving relevant information from external knowledge sources. Recent RAG systems are often built as complex pipelines comprising multiple interconnected modules  (gao2023retrieval, ), including query rewriting ma2023query ; jiang2024rag , first-stage retrieval lewis2020retrieval ; shao2023enhancing , re-ranking salemi2024learning ; salemi2024towards , document preprocessing ke2024bridging ; li2024rag , and answer generation shao2023enhancing ; su2024dragin .
@@
The complexity of RAG systems makes their optimization particularly challenging. Standard supervised fine-tuning (SFT) optimizes each module independently using human-annotated data. However, this often results in misalignment between the objectives of individual components and the overarching goal of the system of generating high-quality results. For example, retrieval modules are frequently trained on human-labeled relevance data to optimize metrics such as nDCGjarvelin2002cumulated . However, this process does not address the disconnect between document relevance and response quality - documents with high relevance scores do not always contribute to generating accurate answers  (cuconasu2024power, ).
@@
To address this issue, existing work on end-to-end optimization for RAG, such as  lee2019latent ; guu2020retrieval ; lewis2020retrieval ; salemi2024towards aims to propagate rewards from the final output to intermediate modules using techniques such as attention distributions izacard2020distilling , generation probability lewis2020retrieval ; zamani2024stochastic , and expectation maximization (EM) iterations singh2021end ; salemi2024learning . However, earlier approaches primarily focus on simplified pipelines with only two components-a retriever and a generator-and fail to provide a generalizable framework for jointly optimizing complex systems with multiple components and richer interdependencies. More recent methods attempt to eliminate the need for module-specific rewards by leveraging algorithms like Direct Preference Optimization (DPO) rafailov2024direct and Proximal Policy Optimization (PPO)schulman2017proximal . Nonetheless, these methods still concentrate on optimizing individual RAG modules in isolation, without adequately modeling the collaborative dynamics between interacting components li2024rag ; ma2023query ; ke2024bridging . Effectively capturing interdependencies among multiple modules and jointly optimizing complex RAG architectures remains an open research challenge.
@@
In this paper, we propose a novel approach called the Multi-Module joint Optimization Algorithm (MMOA-RAG) to enable joint optimization across multiple modules in a RAG system. Our framework treats each intermediate component in the RAG pipeline as an agent and formulates the optimization process as a Cooperative Multi-Agent Reinforcement Learning (Co-MARL) problem, where the agents (i.e., modules) work together to maximize a shared reward for the final outcome.
@@
Specifically, we focus on applying MMOA-RAG to a RAG pipeline that includes four key modules: a query rewriter, a fixed document retriever, a document selector, and an answer generator. Our primary objective is to optimize these modules by defining the final reward as the correctness of the generated response, measured by the F1 score against the ground-truth answer. To achieve this, we leverage the Multi-Agent PPO (MAPPO) algorithm yu2022surprising , which facilitates collaborative optimization in a fully cooperative setting. This means that all modules work in a cooperative way, with their optimization goals aligned toward producing high-quality answers. Unlike previous methods that rely on DPO zhu2024information ; zhu2024atm or PPO ma2023query ; ke2024bridging , MMOA-RAG offers greater flexibility for different pipeline designs and excels at promoting collaboration among multiple modules. This end-to-end optimization ensures that each module’s objectives are consistently aligned with the overarching goal of generating accurate responses.
@@
To demonstrate the effectiveness of the MMOA-RAG modeling and optimization approach, we conducted experiments on three publicly available QA datasets, HotpotQA yang2018hotpotqa , 2WikiMultihopQA ho2020constructing and AmbigQA min2020ambigqa , based on Llama-3-8B-Instruct dubey2024llama . The experimental results indicate that MMOA-RAG achieves better performance than a series of existing optimization methods for RAG. Additionally, we performed extensive ablation studies to investigate the effectiveness and advantages of jointly optimizing mulitple modules in the RAG system and the generalizability of MMOA-RAG across different RAG pipelines.
@@
  • We employ a multi-agent reinforcement learning algorithm to jointly optimize a sophisticated RAG system with four key modules: a query writer, a fixed document retriever, a document selector, and an answer generator.
@@
  • We conduct extensive experiments to verify and demonstrate the effectiveness and generalizability of the proposed framework.
2 Related Works
End-to-end Optimization in OpenQA  Lewis et al. lewis2020retrieval introduces Retrieval-Augmented Generation as RAG, which combines pre-trained language models with non-parametric memory for improved performance on knowledge-intensive NLP tasks. And some other methods lee2019latent ; guu2020retrieval ; izacard2020distilling ; zamani2024stochastic proposed end-to-end optimizing framework for OpenQA system.
RAG without Parameters Updating  Some works jiang2023active ; xu2024search ; su2024dragin ; wang2024astute design a novel RAG mechanism without parameters updating, enhancing the performance of LLMs on question answering tasks.
RAG with Parameters Updating  Some works xu2024unsupervised ; zhao2024longrag ; weiinstructrag optimize RAG with supervised fine-tuning. PPO schulman2017proximal is used by some works schulman2017proximal ; ma2023query ; ke2024bridging ; gao2024smartrag ; jiang2024rag to fine-tune LLMs. Specifically, Search-r1 jin2025search and R1-Searcher song2025r1 both use answer-based reward to improve the reasoning in RAG. Additionally, DPO rafailov2024direct or similar alignment algorithms are used to optimize LLMs in RAG task rafailov2024direct ; zhu2024information ; zhu2024atm ; zhao2024seer ; li2024rag . More detailed related works can be seen in Appendix A.
3 Method
3.1 Modeling RAG as Co-MARL
@@
We define the tuple $\left\langle\mathcal{G},\mathcal{O},\mathcal{A},\mathcal{R}\right\rangle$, where $\mathcal{G}$ denotes the set of agents in the Co-MARL system, $\mathcal{O}$ represents the observation information available to each agent, $\mathcal{A}$ constitutes the action space accessible to each agent, and $\mathcal{R}$ is the reward shared among all agents. The ultimate aim is to maximize this shared reward, thereby achieving higher evaluation metrics and enhancing the overall performance of the RAG system.
In this paper we utilizes Multi-Agent PPO (MAPPO) yu2022surprising , which is an extension of the PPO algorithm schulman2017proximal for multi-agent environments, to optimize the policy for each agent in the Co-MARL framework. In fully cooperative settings, unlike PPO, which focuses on single-agent scenarios with individual reward, MAPPO employs a shared global reward to promote cooperation among all agents.
3.2 Overall of MMOA-RAG
Figure 1: The overall framework of MMOA-RAG.
@@ 3.2 Overall of MMOA-RAG
  • Query Rewriter reformulates the initial question $q$, which may be too complex or ambiguous to resolve with a single retrieval, into a set of sub-questions denoted as subqsubq.
@@
Since the Query Rewriter, Selector, and Generator modules can all be implemented using LLMs, they can be treated as RL agents ouyang2022training , enabling parameter updates through reward signals. To optimize computational efficiency, these three modules can share the same LLM. Additionally, given the difficulty of modeling the Retriever module as an RL agent, we use a fixed Retriever and regard it as a part of the environment  [^2] Recent studies in generative IR (see (li2024matching, ) for a survey) have explored using generative models for retrieval. But we choose a more traditional dense retrieval model izacard2021unsupervised as the first-stage retriever and leave the optimization of the first-stage retriever for future work..
@@
The focus of the MMOA-RAG framework is on the collaborative optimization of multiple modules to align their individual objectives with the ultimate goal of generating high-quality answers. We use metrics from the Generator’s predicted answer $Ans_{\text{predict}}$, such as F1 score, as a shared reward $R_{\text{shared}}$. Given the fully cooperative nature of the modules in the RAG system, $R_{\text{shared}}$ can be used to train all agents, a common approach in existing MARL literature QMIX ; yu2022surprising ; PTDE . Additionally, to ensure training stability and accelerate convergence in the multi-agent system, we design penalty terms $P_{QR}$, $P_{S}$, and $P_{G}$ for each agent. A more detailed explanation will be provided in Section 3.3.
3.3 Detailed Configuration for Each Agent
In this section, we will provide a detailed explanation of each element in the tuple $\left\langle\mathcal{G},\mathcal{O},\mathcal{A},\mathcal{R}\right\rangle$ mentioned in Section 3.1. Here, $\mathcal{G}=\{\text{Query Rewriter (QR), Selector (S), Generator (G)}\}$ represents all agents. In the following, we introduce the essential elements for each agent $i\in\mathcal{G}$: the observation information $O_{i}\in\mathcal{O}$, the action space $A_{i}\in\mathcal{A}$, and the reward function $R_{i}$.
3.3.1 Elements of Query Rewriter
Observation of Query Rewriter is defined as Equation (1), which contains prompt of Query Rewriter $Prompt_{QR}$ (as shown in Table 3) and the initial question $q$.
$O_{QR}=\left\{Prompt_{QR},q\right\}$ (1)
Action Space of Query Rewriter corresponds to the vocabulary of LLMs, denoted as $\mathcal{V}$, as we prompt the LLM to generate one or more sub-questions based on $q$.
$A_{QR}=\mathcal{V}$ (2)
$R_{QR}=R_{\text{shared}}+P_{QR}$ (3)
3.3.2 Elements of Selector
Observation of Selector is defined as Equation (4), which contains prompt of Selector $Prompt_{S}$ (as shown in Table 4), the initial question $q$ and the candidate documents set $D$ with $K$ documents.
$O_{S}=\left\{Prompt_{S},q,D\right\}$ (4)
$A_{S}=\left\{\text{"{0}"},\text{"{1}"},\ldots,\text{"{K-1}"},\text{"{Document}"},\text{"{,}"}\right\}$ (5)
$R_{S}=R_{\text{shared}}+P_{S}$ (6)
3.3.3 Elements of Generator
Observation of Generator is in Equation (7), containing prompt of Generator $Prompt_{G}$ (as shown in Table 5), the initial question $q$ and the selected candidate documents set $D_{\text{selected}}$ given by Selector.
$O_{G}=\left\{Prompt_{G},q,D_{\text{selected}}\right\}$ (7)
Action Space of Generator $A_{G}$ is the same as Query Rewriter.
$A_{G}=A_{QR}=\mathcal{V}$ (8)
Reward Function of Generator contains $R_{\text{shared}}$ and penalty term $P_{G}$, which is used to constrain the model from generating excessively long content. When the generated answer exceeds a certain length, $P_{G}$ is set to -0.5; otherwise, it is set to 0. In fact, the values of each penalty $P_{i}\ (i\in\mathcal{G})$ are mostly 0, and they only become negative when the output does not meet the requirements.
$R_{G}=R_{\text{shared}}+P_{G}$ (9)
3.4 Training Process of MMOA-RAG
3.4.1 Warm Start with SFT
In preparation for joint optimization of multiple modules using Multi-Agent PPO, it is essential to perform a warm start for each trainable module. The warm start enables the model to better adhere to instructions across diverse tasks and reduces the exploration space during MARL joint training, thereby enhancing the efficiency of exploration and exploitation.
Within the MMOA-RAG framework, there are three trainable modules: the Query Rewriter, the Selector, and the Generator. Consequently, we construct the training data for the SFT of each corresponding task and perform the SFT to get the warm-up checkpoints for each trainable modules. The details of constructing training data can be seen in Appendix B.
3.4.2 Multi-Agent Optimization